OpenStudy (anonymous):

Choose the correct simplification of (4a2 + 9a − 5) + (6a2 − 5). 10a2 + 9a 10a2 + 9a − 10 −2a2 + 9a −2a2 + 9a − 10

OpenStudy (anonymous):

4a2+4a2+9a+(-5)+(-5) =?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

10a2 + 9a − 10

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

u add all equal terms

OpenStudy (anonymous):

please dont be vague. explain how to get the answer. tell me how to solve (4a2 + 9a − 5) + (6a2 − 5)

OpenStudy (anonymous):

@anas2000 please explain

OpenStudy (anonymous):

Group the same terms. Whatever has a2 is added together, those with a are added together and the constants are added together.

OpenStudy (anonymous):

@mikaa_toxica13

OpenStudy (raden_zaikaria):

4a2 + 9a − 5+ 6a2 − 5

OpenStudy (raden_zaikaria):

4a2 + 6a2 = 10 a2 and then − 5− 5 = -10 and then 9a

OpenStudy (raden_zaikaria):

10 a2 + 9a -10 this answer!!